what is the maximum possible concentration of Ni2+ ions in a solution containing 0.15 M HCl and 0.1 M H2S. given that Ksp(NiS)= 2*10^-21 and [S2-]=4*10^-21.

Grade:11

1 Answers

Aman Bansal
12 years ago

Dear Parth,

find the Ka for H2S acid as l.l x l0-7 in pure water.

Ka expression l.l x l0-7 = (H+)^2(S=) over (H2S)

but in this solution the pH is 3.4, so next find the H+ concentration.

H+) – antilog of -pH or antilog of -3.4 which equals 3.98 x l0-4 moles per liter (H+) conc.

Plug this value into your Ka expression for the weak acid equilibrium and solve for maximum concentration allowed for the Sulfide ion.

Ka = l.l x l0-7 = (3.98 x l0-4)(S=) over .13M H2S)

l.43 x l0-8 = 3.98 x l0-4(S=)

so (S=) = 3.59 x l0-5 allowable S= concentration that can combine with Ni ion.

Now plug this number into the Ksp expression for NiS

Ksp = l.3 x l0-25 = (Ni++)(S=) = (Ni++)(3.98 x l0-5)

so (Ni++) = l.3 x l0-25 divided by 3.98 x l0-5 which equals 3.62 x l0-21st moles/liter Ni++ max concentration

Best Of luck

Plz Approve the answer...!!!!

Cracking IIT just got more exciting,It’s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and winexciting gifts by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple click here to download the toolbar….
So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert